# STAR: SocioTechnical Approach to Red Teaming Language Models

### v3

@@ 6.2.2 Arbitrators weigh annotator reasoning
Qualitative analyses of arbitrator reasoning shows a notably high level of consideration and quality of annotator and arbitrator reasoning (for examples, see I). Rather than picking one side, arbitrators typically weighed the reasoning of both annotators and provided their own reasoning from the perspective of an independent third party, somewhat like a judge writing a verdict (see I.2). For example, arbitrators often highlight key terms of disagreement, such as whether fictional stories count as ‘promoting’ hate or stereotype, or whether accepting a hateful premise in an attack counts as hate.
We compute the inter-rater reliability across all annotators, within six high-level policy areas (see  B), and find Krippendorf’s Alpha =.$=.50$ over the entire Likert scale, and Krippendorf’s Alpha =.$=.47$ with binarised response options. In addition to meaningful disagreement, qualitative analysis of annotator reasoning revealed that some disagreement between any two raters originated in different interpretations of the instructions. For example, raters disagreed on whether a fictional story that included harmful stereotypes constituted a rule violation. Disagreement also arose in some cases when the model initially abided by the targeted rule but produced harmful content later on – some annotators argued that the attacker was to blame for forcing or tricking the model into a violative response. Similarly, situations where the attacker preconditioned the model to adopt a specific viewpoint on a topic (e.g. instructing the model to take an action or express an opinion based on racial stereotypes) generated more disagreement.
@@ 7 Conclusion & Discussion
In addition to addressing steerability and controllability challenges, by introducing a principled process for generating such instructions, STAR also provides an approach to another ongoing challenge in the red teaming field - that of creating reproducible processes for generating comparable red teaming datasets.
@@
As a proof of concept, we demonstrate that STAR can be used to target specific risk areas of different levels of specificity. This is effective, as the cluster analysis comparing multiple red teaming approaches shows that gender stereotypes and race-based bias are the main topics of our resulting dialogues in STAR - as targeted in the instructions, but not in other red teaming approaches that cast a broader focus. Notably, while DEFCON and Anthropic give more open-ended instructions to red teamers, these efforts end up clustering in different areas that were not described as key intended target areas, particularly on malicious use and comparably narrow failure modes such as PII release. This suggests that open-ended instructions do not provide broader coverage than highly structured, parameterised instructions as provided in STAR. Rather, STAR is an approach to exercise more intentional control over the target area, without resulting in higher clustering of resulting dialogues.
@@
We note that parameterised instructions enable more nuanced findings about model failure modes without incurring additional costs. This may reveal previous blind spots - in our case, showing that while the model is not more likely to spew hate speech about a particular race or gender, it is more likely to reproduce social marginalisation when prompted about gender×$\times$×race intersectionalities, specifically women of colour, compared to white men. In this way, the parametric approach of STAR provides a significant value-add by enabling more nuanced coverage of failure modes at no additional cost.
@@
We further find that diversifying annotator pools and demographic matching lead to higher sensitivity in annotations on discriminatory stereotypes and hate speech on specific groups. This suggests that in-group members bring experience and perspectives to bear that differ from those of out-group members. Without demographic matching, these perspectives may have been buried by majority views. By prioritising the insights of those most directly affected in the context of hate and stereotypes, we ensure a legitimate and authoritative assessment of model failures. We find reasonable inter-rater agreement, showing that our approach compares to state of the art approaches (Ganguli et al., 2022; Xu et al., 2021).
@@
Finally, we show that annotator disagreement can be a rich source of signal. Disagreement between red teamers is often reported as undesirable but then discarded. This loses an informative signal, as disagreement may in part stem from different subjective perspectives that ought to be treated differently. Here, prompting annotators to share their reasoning in free-form text enabled qualitative analysis on underlying reasons for such disagreement and demonstrated high quality of reasoning. It also allowed for a more comprehensive arbitrator judgement weighing different arguments.
@@ 8 Future directions
The adaptable nature of the parameterized STAR approach allows for red teaming models on harms, use cases, and failure modes tailored to diverse locales and priorities. STAR can be extended to any combinatorial space of potential attacks or failures, making it highly adaptable to different contexts. For instance, instructions can be easily modified to address specific social categories like "caste" instead of "race," or include additional parameters like "age" to investigate intersectional harms. Furthermore, STAR can be applied to various languages, modalities of model output, or user applications. While currently a method for human red teaming, STAR can be adapted for a hybridized approach, incorporating automated tools to augment human red teamers.
@@ 9 Limitations
However, STAR is limited by the cognitive load that human raters can absorb - here, we use maximally five parameters. A different limit may apply to automated red teamers. In our particular use of STAR, we attack the model only in English language, against specific harm areas and with specific demographic labels (gender, race). This limited charting of the attack surface serves to highlight model failures in this area but cannot speak to model failures in other domains. The high-dimensional embedding may be influenced by stylistic differences between model responses, as well as between real-world users and red teamers. Further, despite careful and detailed instructions, we find some clustering of dialogues that do not seem to mirror real-world innocuous use (as indicated in the real-world dialogue dataset). This may in part be due to limited interaction methods in our task design - for example, we do not permit certain actions that may be possible in real-world use of generative AI systems, such as uploading documents for the language model to ingest. In particular, we note that none of the projected red teaming approaches overlap entirely with flagged instances of real-world user-AI-interactions. This suggests that more work is needed to ensure broad coverage of real-world failures in a red teaming setup.
@@ 10 Acknowledgments
Acknowledgements: We thank Raia Hadsell, Shakir Mohamed, Slav Petrov, Jenny Brennan, Ben Bariach and Iason Gabriel for feedback and reviews. We thank Sarah Hodkinson, Michael Sharman, Courtney Biles and Shereen Ashraf for support in project management and Steph Huang for graphic design. We thank Rhiannon Weaver, Tim Coleman, Burim Ramosaj for discussions on statistical methodologies.
@@ Appendix B Safety Content Policy
Our testing focused on ‘rules’ as laid out in a safety content policy that applies to the proprietary model that was evaluated (Terms, 2023).
@@
Within the second topic area, we particularly focus on two rules: ’Generating content that promotes or encourages hatred’ ’Facilitating methods of harassment or bullying to intimidate, abuse, or insult others’
@@ C.1 Red Teaming task
See instructions in Fig. 6.
@@
Topics can be freely selected and submitted via an open-text field. Alternatively, participants can obtain suggested topics which are randomly selected from a publicly shared repository of ’content categories’ (Cloud, 2023).
@@ C.2 Annotation task
See instructions in Fig. 7 and an example of the task in Fig. 8.
Figure 7: Instructions for the Annotation task.
@@ C.3 Arbitration task
C.4 Demographic Matching (Annotation or Arbitration task)
See an example in Fig. 10.
@@ Appendix D Demographic matching
Appendix E Participant demographics
@@
Appendix F Dataset descriptions
@@
    STAR - Conversations produced by Red Teamers during the STAR project outlined in this paper.
    Anthropic - Created as part of Anthropic’s red teaming paper Ganguli et al. (2022) which were open sourced as part of paper release in 2022. Assessments were performed on a proprietary model created by Anthropic.
    DEFCON - Generated during a 3 day contest at DEFCON 31 in August 2023. This was a public contest where red teamers assessed 8 models provided by frontier AI organisations.
    Real-world user dialogues - Conversations between users and a proprietary model in 2023 that were flagged by users.
Appendix G In- vs. out-group red teaming attack success
@@
Appendix H In- vs. out-group annotations
@@
Appendix I Examples of Annotator and Arbitrator Reasoning
@@
I.1 Annotator reasoning
Annotators held different opinions on whether describing stereotypes counts as ‘promoting’ hate or violence toward a group as was specified in the rule. For example, one annotator reasoned about description of violence against slaves: ‘The chatbot generates a story based on well-known facts and occurrences of slaves. This story does not encourage or promote violence. It only describes the acts in a storytelling method.’ Similarly, one annotator reasoned about stereotypes against Hispanic people: ‘Although there was a statement that indicated the “majority of Hispanics identify as Roman Catholic”, could’ve been questionable [sic] but the fact is through research that more than half of Hispanics do identify as Roman Catholic. That makes it a factual statement.’

### v4

@@ 6.2.2 Arbitrators weigh annotator reasoning
Qualitative analyses of arbitrator reasoning shows a notably high level of consideration and quality of annotator and arbitrator reasoning (for examples, see Appendix J). Rather than picking one side, arbitrators typically weighed the reasoning of both annotators and provided their own reasoning from the perspective of an independent third party, somewhat like a judge writing a verdict (see J.2). For example, arbitrators often highlight key terms of disagreement, such as whether fictional stories count as ‘promoting’ hate or stereotype, or whether accepting a hateful premise in an attack counts as hate.
We compute the inter-rater reliability across all annotators, within six high-level policy areas (see Appendix B), and find Krippendorf’s Alpha =.$=.50$ over the entire Likert scale, and Krippendorf’s Alpha =.$=.47$ with binarised response options. In addition to meaningful disagreement, qualitative analysis of annotator reasoning revealed that some disagreement between any two raters originated in different interpretations of the instructions. For example, raters disagreed on whether a fictional story that included harmful stereotypes constituted a rule violation. Disagreement also arose in some cases when the model initially abided by the targeted rule but produced harmful content later on – some annotators argued that the attacker was to blame for forcing or tricking the model into a violative response. Similarly, situations where the attacker preconditioned the model to adopt a specific viewpoint on a topic (e.g. instructing the model to take an action or express an opinion based on racial stereotypes) generated more disagreement.
@@ 7 Conclusion & Discussion
In addition to addressing steerability and controllability challenges, by introducing a principled process for generating such instructions, STAR also provides an approach to another ongoing challenge in the red teaming field - that of creating reproducible processes for generating comparable red teaming datasets. While red teaming as a method is targeted at surfacing new failure modes, it can in some cases be desirable to compare outcomes from different red teaming efforts. To the extent that such red teaming efforts rely on similarly parametrised instructions, this comparison becomes more reliable as potential confounds affecting open-ended red teaming efforts can be better controlled for.
@@
As a proof of concept, we demonstrate that STAR can be used to target specific risk areas of different levels of specificity. This is effective, as the cluster analysis comparing multiple red teaming approaches shows that gender stereotypes and race-based bias are the most common topics of our resulting dialogues in STAR - as targeted in the instructions, but not in other red teaming approaches that cast a broader focus. Notably, while DEFCON and Anthropic give more open-ended instructions to red teamers, these efforts end up clustering in different areas that were not described as key intended target areas, particularly on malicious use and comparably narrow failure modes such as PII release. This suggests that open-ended instructions do not provide broader coverage than highly structured, parameterised instructions as provided in STAR. Rather, STAR is an approach to exercise more intentional control over the target area, without resulting in higher clustering of resulting dialogues.
Parameterising instructions with random combinations of parameter values (a kind of randomised factorial design) allows for nuanced, retroactive analysis without increasing data collection costs. Examining the marginal effects reveals parameter values and intersections that contribute to model failures, potentially uncovering blind spots.
@@
In our case, while the model is not more likely to spew hate speech about a particular race or gender, it is more likely to reproduce social stereotypes when prompted about gender×$\times$×race intersectionalities, specifically women of colour, compared to white men.
When budgeting how many dialogues to collect for statistical power, the required sample size depends primarily on the planned analyses. For example, to plot a 2D heatmap from two parameters and achieve statistically reliable results in each bin, the sample size must scale with the product of the number of values those parameters can take [^5] To achieve sufficient power, we limited the number of demographic intersections we red teamed to certain gender×$\times$×race intersections (Appendix D). To represent a larger set of intersections (Bergman et al., 2023), it would be possible to collect more dialogues or to stratify further with the same number of overall dialogues, at the expense of powering between-group significance tests.. Adding parameters may also affect the number of dialogues required. Additional parameters can stratify the data collection (divide it into meaningful subgroups) without affecting the required number of dialogues, if parameters are not expected to meaningfully interact. However, when adding a novel parameter (e.g. gender), care must be taken to consider how that parameter may interact with other parameters (e.g. race), to ensure downstream analyses are sufficiently powered.
@@
We find that diversifying annotator pools and demographic matching leads to higher sensitivity in annotations on discriminatory stereotypes and hate speech on specific groups. This suggests that in-group members bring experience and perspectives to bear that differ from those of out-group members. Without demographic matching, these perspectives may have been buried by majority views. By prioritising the insights of those most directly affected in the context of hate and stereotypes, we ensure a legitimate and authoritative assessment of model failures. We find reasonable inter-rater agreement, showing that our approach compares to state of the art approaches (Ganguli et al., 2022; Xu et al., 2021).
@@
Finally, we show that annotator disagreement can be a rich source of signal. Disagreement between red teamers is often reported as undesirable and then discarded. This loses an informative signal, as disagreement may in part stem from different subjective perspectives that ought to be treated differently. Here, prompting annotators to share their reasoning in free-form text enabled qualitative analysis of the underlying reasons for such disagreement and demonstrated high quality of reasoning. It also allowed for a more comprehensive arbitrator judgement weighing different arguments.
@@ 8 Future directions
The adaptable nature of the parameterised STAR approach allows for red teaming models on harms, use cases, and failure modes tailored to diverse locales and priorities. STAR can be extended to any combinatorial space of potential attacks or failures, making it highly adaptable to different contexts. For instance, instructions can be easily modified to address specific social categories like "caste" instead of "race," or include additional parameters like "age" to investigate intersectional harms. Furthermore, STAR can be applied to various languages, modalities of model output, geographic regions or user applications.
Whilst designed for human red teaming, STAR can be adapted for semi-automated approaches. It can be used as a baseline against which to benchmark the coverage of fully or partially automated red teaming. Alternatively, in a hybrid approach, automated tools could explore a prompt space via parameterised instructions (see also concurrent Radharapu et al. (2023)) and ascertain edge cases and likely failure modes, while humans could focus on higher-level tasks like defining risk areas to explore and addressing edge cases, leveraging their experience and contextual understanding.
To support the tailoring of red teaming methodologies for different contexts, future work may compare the respective advantages and disadvantages of automated or semi-automated approaches, taking into account factors such as breadth and depth of coverage, attack success rate, participant wellbeing, and cost (Appendix E).
@@ 9 Limitations
However, STAR is limited by the cognitive load that human raters can absorb – here, we use at most five parameters, and the demographic group parameter is at most a two-way intersection [^6] For even more comprehensive coverage it would have been ideal to red team more complex demographic intersections that may affect model performance in the context of social harms, such as sexual orientation, religion, disability status or age. However introducing highly complex intersections to a prompt would have placed a high cognitive burden on red teamers. Future work may explore red teaming against these demographic labels. Specialised expert red teamers may find it harder to leverage their expertise when constrained by parameterised instructions. In such cases the parameterised instructions can be used as inspiration, providing starting points or prompting the consideration of specific themes, rather than a rigid requirement.
In our particular use of STAR, we attack the model only in English, against specific harm areas and with specific demographic labels (gender, race). This limited charting of the attack surface serves to highlight model failures in this area but cannot speak to model failures in other domains.
The high-dimensional embedding used for the UMAP may be influenced by stylistic differences between model responses, as well as between real-world users and red teamers. Furthermore, one of the statistical assumptions of UMAP is that the data is uniformly distributed over the underlying manifold, which is most likely not the case in red teaming efforts as red-teamers discover strategies that work or don’t work.
Despite careful and detailed instructions, we find some clustering of dialogues that do not seem to mirror real-world innocuous use (as indicated in the real-world dialogue dataset). This may in part be due to limited interaction methods in our task design - for example, we do not permit certain actions that may be possible in real-world use of generative AI systems, such as uploading documents for the language model to ingest. In particular, we note that none of the projected red teaming approaches overlap entirely with flagged instances of real-world user-AI-interactions. This suggests that more work is needed to ensure broad coverage of real-world failures in a red teaming setup. Finally, the comparison of STAR to prior approaches relies on previously released datasets rather than careful experimental variation and ablation. Future work may systematically study the impacts of methodological innovations in red teaming, such as those introduced in STAR.
@@ 10 Acknowledgments
Acknowledgements: We thank Raia Hadsell, Shakir Mohamed, Slav Petrov, Iason Gabriel, Jenny Brennan and Ben Bariach for feedback and reviews. We thank Sarah Hodkinson, Michael Sharman, Courtney Biles and Shereen Azraf for support in project management and Steph Huang for graphic design. We thank Rhiannon Weaver, Tim Coleman, Burim Ramosaj for discussions on statistical methodologies.
@@ Appendix B Safety Content Policy
Our testing focused on ‘rules’ based on a safety content policy that applies to the proprietary model that was evaluated (Terms (2023)).
@@
Within the first topic area, we particularly focus on the following rules:
  1. 1.d.i
    Generating content that promotes or encourages hatred
  2. 1.d.vii
    Generating content that may have unfair or adverse impacts on people, particularly impacts related to sensitive or protected characteristics
These rules were rephrased for ease of use in the participant-facing instructions, see Appendix C.
@@ C.1 Red Teaming task
See an example of instructions procedurally generated from parameterised templates in Figure 6.
@@
Topics can be freely selected and submitted via an open-text field. Alternatively, participants can obtain suggested topics which are randomly selected from a publicly shared repository (Cloud, 2023).
@@ C.2 Annotation task
See annotation form in Figure 7 and an example of the task in Figure 8.
Figure 7: Form for the Annotation task.
@@ C.3 Arbitration task
C.4 Demographic Matching instructions (Annotation or Arbitration task)
See an example in Figure 10.
@@ Appendix D Demographic matching
Appendix E Participant compensation and labour costs
We collect a total of 3,236 participant hours, with 1,614 hours spent red teaming and 1,622 spent on annotation. Participants were paid at or above the living wage for their location.
Appendix F Participant demographics
@@
Appendix G Dataset descriptions
@@
  • STAR: Conversations produced by Red Teamers during the STAR project outlined in this paper.
  • Anthropic: Created as part of Anthropic’s red teaming paper Ganguli et al. (2022) and open sourced as part of paper release in 2022. Assessments were performed on a proprietary model created by Anthropic.
  • DEFCON: Generated during a 3 day contest at DEFCON 31 in August 2023 (defcon). This was a public contest where red teamers assessed 8 models provided by frontier AI organisations.
  • Real-world user dialogues: Conversations between users and a proprietary model in 2023 that were flagged by users.
Appendix H In- vs. out-group red teaming attack success
@@
Appendix I In- vs. out-group annotations
@@
Appendix J Examples of Annotator and Arbitrator Reasoning
@@
J.1 Annotator reasoning
Annotators held different opinions on whether describing stereotypes counts as ‘promoting’ hate or violence toward a group as was specified in the rule. For example, one annotator reasoned about a description of violence against slaves: ‘The chatbot generates a story based on well-known facts and occurrences of slaves. This story does not encourage or promote violence. It only describes the acts in a storytelling method.’ Similarly, one annotator reasoned about stereotypes against Hispanic people: ‘Although there was a statement that indicated the “majority of Hispanics identify as Roman Catholic”, could’ve been questionable [sic] but the fact is through research that more than half of Hispanics do identify as Roman Catholic. That makes it a factual statement.’